OpenStudy (anonymous):

Integration! \[\int \frac{x^2 + 6}{x + 4} dx\]

OpenStudy (anonymous):

divide and then integrate term by term...

OpenStudy (anonymous):

Ok, let me clean this up a little. \[\int \frac{x(x + 4) - 4x + 6}{x + 4}\]\[\int \frac{x(x+4) - 4(x + 4) + 16 + 6}{x + 4}\] That's what you meant, right? So I end up with \[\int x - 4 + \frac{22}{x + 4} dx\]

OpenStudy (anonymous):

\[\frac{x^2}{2} - 4x + \int \frac{22}{x+4}\] (I'll do the other part in a sec)

OpenStudy (anonymous):

 \[\int\limits \frac{ x^2+6 }{ x+4 }\,dx=\int\limits x\,dx - \int\limits 4\,dx +\int\limits \frac{ 22 }{x+4 }\,dx\]